OpenStudy (anonymous):

what is 10/-7 as a mixed number

OpenStudy (anonymous):

1 3/10

OpenStudy (anonymous):

can you help me find the answer to this equation? -12= -7-7/2s

OpenStudy (anonymous):

\[12=-7-7/2s\]

OpenStudy (anonymous):

7/2s is a fraction right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, so first of, lets get the easy done, lets add the -7 over to both sides.

OpenStudy (anonymous):

ok -5

OpenStudy (anonymous):

-5=7/2s

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Lets multiply by 2s to get s alone

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what would the final answer be then? I'm really bad at equations

OpenStudy (anonymous):

s=-7/10

OpenStudy (anonymous):

now, once you got your answer, plug it in to ur original equation and see if it works out, if it does, you know you got the question correct, if not, its problably incorrect.

OpenStudy (anonymous):

wouldn't it be -10/7?

OpenStudy (anonymous):

Try it out :D

OpenStudy (anonymous):

I redid the equation and I got -10/7

OpenStudy (anonymous):

Now I need to turn that into a mixed number

OpenStudy (anonymous):

Try this site: http://library.thinkquest.org/J002328F/impmix.htm

OpenStudy (anonymous):

oh so the answer is 1 3/7